OpenStudy (anonymous):

Is it possible to design a tank such that the depth of the liquid, h, will be a linear function of time, t? question related to Torricelli's law

OpenStudy (irishboy123):

yes

OpenStudy (anonymous):

haha ok , plz elaborate

ganeshie8 (ganeshie8):

\(h = h_0 + mt\) should do right ?

ganeshie8 (ganeshie8):

Oh you want to design a tank that honors a linear function, interesting !

ganeshie8 (ganeshie8):

\(p + \frac{1}{2} \rho v^2 + \rho g y\) = constant

ganeshie8 (ganeshie8):



OpenStudy (alexandervonhumboldt2):

wow

OpenStudy (alexandervonhumboldt2):

ROFL

ganeshie8 (ganeshie8):

Taking \(y=0\) at the bottom and noticing that both the surfaces are exposed to atmosphere, do we get : \[p_0+\frac{1}{2}\rho v_0^2+\rho g h = p_0+\frac{1}{2}\rho v^2+\rho g (0) \] where \(v_0\) = speed at which top surface is moving \(v\) = speed of leaking water

ganeshie8 (ganeshie8):

That simplifies to \[v^2- v_0^2= 2g h \]

OpenStudy (irishboy123):

you need to include the diameter of the "hole in the bottom" as a parameter after that, it is Bernoulli and continuity equations

ganeshie8 (ganeshie8):

pressure only depends on height, so it seems the geometry of tank is of no importance ?

OpenStudy (anonymous):

Alas, the rate of height change is dependent on volume so it appears it does depend on shape (unless i've misinterpreted the question)

ganeshie8 (ganeshie8):

Oh, are you saying we can adjust the diameter of the hole and make the leaking rate somehow linear ?

OpenStudy (irishboy123):

but the size of the hole matters?

OpenStudy (anonymous):

@IrishBoy123 - yes hole size matters in a non-ideal system, but I think head loss is overkill for this problem...

OpenStudy (anonymous):

The change in the depth of liquid is related to the volumetric flow out of the tank. Assuming the hole doesn't change shape, the volumetric flow is related directly to the velocity of the flow. If we use Toricelli's law, we know that v=sqrt(2gz) or that v is proportional to the square root of the height. If you vary the cross section such that the cross sectional area changes with the height then i think it would be possible in an ideal system

OpenStudy (anonymous):

Though please correct me if I've made any egregious assumptions or mistakes

OpenStudy (irishboy123):

@ LifeEngineer yep!! but look at the OP's question it is pretty meaningless unless you are doing other things like keeping the volume constant.

ganeshie8 (ganeshie8):

I think we need to have \(v\propto \sqrt{t}\)... does below geometry work ? not so sure..